OpenStudy (anonymous):

What is the equation 5x – 2y = 10 in slope-intercept form? What is the slope and y-intercept?

OpenStudy (anonymous):

Slope Intercept from is y=mx+b.. where m is the slope and b is the y-intercept... So solve for y subtract 5x to both sides -2x=-5x+10.... divide by -2 y= 5/2x -5 So the slope is 5/2 and the y intercept is -5

OpenStudy (okra77):

slope is 5/2 equation is y=5/2x-5